OpenStudy (anonymous):

Compute the indicated derivative using the chain rule x = 1 − t/4, y = 7t − 1 ; dy/dx dy/dx : ____

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ dy }{ dt }*\frac{ dt }{ dx }\]

OpenStudy (ksaimouli):